Status Report

NOAA SATOPS Morning Report: Wednesday, June 16, 2010

By SpaceRef Editor
June 16, 2010
Filed under , , ,

Customers or users who want to call to ask about problems or report outages can use the following contacts:

– For questions about data quality (noisy data, dropouts, schedules), GOES or POES products, or status of spacecraft can call the SAB Shift Supervisor at 301-763-8444

– To report an outage or document a problem, call the ESPC Operations Crew Lead at 301-457-5218. All LRIT data users are encouraged to report problems to: LRIT@noaa.gov

DOY 167

GOES
All GOES operations were nominal over the past 24 hours.
No anomalies affecting product delivery or s/c health & safety.
For GOES status and other information please see:
http://www.oso.noaa.gov/goes

POES
N15 NOTE: On DOY 070 N15 Gyro-2 was turned off due to current spikes which could have also impacted the last operational gyro (gyro-1). As a result of the transition to a reduced gyro attitude control mode users may experience significant geolocation errors in their products.
Engineers are reporting a 2.5 degree average attitude error about the yaw axis.

N17 NOTE: Due to a NOAA-17 AVHRR Scan Motor problems AVHRR imagery will experience periods of instability. This cycling of instability is a end of life condition and not expected to permanently improve.

All other POES operations were nominal over the past 24 hours.
No anomalies affecting product delivery or s/c health & safety.
For a detailed status and description of POES see:
http://www.oso.noaa.gov/poes

DMSP
All DMSP operations were nominal over the past 24 hours.
No anomalies affecting product delivery or s/c health & safety.
For more DMSP Information please see:
http://dmsp.ngdc.noaa.gov/dmsp.html

JASON-2
All JASON-2 operations were nominal over the past 24 hours.
No anomalies affecting product delivery or s/c health & safety.

Questions or comments to Ron.Mahmot@noaa.gov

SpaceRef staff editor.